OpenStudy (anonymous):

the ratio of the diameter of the conical tent to its height is 3 to 5.if the height is 10ft. find the amount of air inside

OpenStudy (anonymous):

\[\frac{ d }{ h }=\frac{ 2r }{ h }=\frac{ 3 }{ 5 }\] \[\frac{ 2r }{ 10 }=\frac{ 3 }{ 5 }\] find r calculate volume

OpenStudy (anonymous):

how?